OpenStudy (anonymous):

15/x+4 = x-4/x [Solve this equation.]

OpenStudy (anonymous):

Could you please show me step by step?

OpenStudy (campbell_st):

start by cross multiplying this will remove the fractions.... \[\frac{15}{x + 4} = \frac{x - 4}{x} = 15 \times x = (x + 4) \times (x - 4)\] now you should be able to simplify the problem.

OpenStudy (anonymous):

Is it 15x=x^2-4x+4x-16?

OpenStudy (anonymous):

I don't understand.

OpenStudy (campbell_st):

ok... so you're correct with the expansion. so 15x = x^2 - 16 now put all the terms on one side x^2 - 15x - 16 = 0 now you have a quadratic that can be solved by factorising...

OpenStudy (anonymous):

oh okay. I get it now ^^. thank you..

OpenStudy (anonymous):

I thought I might have been doing it wrong.

OpenStudy (campbell_st):

nope... it all makes sense... so you'll have 2 solutions.